OpenStudy (anonymous):

Suppose f(x)=[sin(x^2-4)]^-1. Identify any points of discontinuity, and determine (giving reasons) if they are removable, infinite (essential), or jump discontinuities.

OpenStudy (anonymous):

is this \[\frac{1}{\sin(x^2-4)}\]/

OpenStudy (anonymous):

i would be undefined a whole raft of places, since sine is periodic

OpenStudy (anonymous):

since \(\sin(0)=0\) it would undefined is \[x^2-4=0\] in other words if \(x=2,x=-2\) but that only gives two values

OpenStudy (anonymous):

also undefined is \(x^2-4=\pi\) etc

OpenStudy (anonymous):

um satellite can you help me with a question after this?

darn it he left ):